OpenStudy (tmostiller):

What is the value of y^-5/x^-3 for x=2 and y=-4? A. 10/3 b. 128 C. -1/128 d. -128. I suck at Algebra

OpenStudy (mathstudent55):

Replace y with the given value and x with the given value. Do you know how to work with negative exponents?

OpenStudy (tmostiller):

Completely lost

OpenStudy (brookegallagher):

When you have negative exponents, you want to switch the places of the variables from numerator to denominator (or vice versa) and make the exponent positive. For example, if you have \[x ^{-y}\] you would make it \[\frac{ 1 }{ x ^{y} }\] For this problem, you'd take the expression and rewrite it as: \[\frac{ x ^{3} }{ y ^{5} }\] Now we can plug in the values for the variables like so: \[\frac{ 2^{3} }{ (-4)^{5} }\] Now solve! \[\frac{ 8 }{ -1024 }\] And simplify! \[\frac{ -1 }{ 128 }\] The answer is C. -1/128. Hope this helps!